OpenStudy (christos):

Linear Algebra, So I have to prove this: "Let a be a 1*n matrix and B an n*p matrix. Show that the matrix product aB can be written as a linear combination of the rows of B, where the coefficients are the entries of a" Can you basically explain to me the problem? I don't want the solution just explain to me ? Thanks

OpenStudy (christos):

@agent0smith

OpenStudy (christos):

@amistre64

OpenStudy (christos):

@Callisto

OpenStudy (christos):

@dan815

OpenStudy (christos):

@dumbcow

OpenStudy (christos):

@ganeshie8

OpenStudy (christos):

@hartnn

OpenStudy (christos):

@hba

OpenStudy (christos):

@phi

OpenStudy (christos):

basically I don't understand what I need to "show"

OpenStudy (amistre64):

Let a be a 1*n matrix A is a 1 row, n column matrix ... and B an n*p matrix. B has n rows and p columns Show that the matrix product aB can be written as a linear combination of the rows of B a linear combination is a "sum" of terms; can you better define what aB operation is? where the coefficients are the entries of a

OpenStudy (amistre64):

im thinking AB is an inner procduct; so you will most likely have to employ the defintion of an inner product

OpenStudy (christos):

hmm does that have anything to do with \[\sum\limits_{}^{} {something} \]

OpenStudy (christos):

this type of thing ?

OpenStudy (amistre64):

a linear combination of the vector <3,2,1> is 3i + 2j + k

OpenStudy (christos):

but here I don't play with vectors or do I ?

OpenStudy (christos):

I think I play with matrices

OpenStudy (christos):

or whatnot

OpenStudy (amistre64):

matrixes are built out of vectors

OpenStudy (christos):

so for example a 5*5 matrice has 25 vectors ?

OpenStudy (amistre64):

each column is a vector from a given span

OpenStudy (amistre64):

a 5x5 has 5 column vectors

OpenStudy (amistre64):

each column vector of a 5x5 matrix represents a spacial vector with 5 parts:\[\begin{pmatrix}x_1\\x_2\\x_3\\x_4\\x_5\end{pmatrix}\]

OpenStudy (christos):

and 5 row vectors ?

OpenStudy (christos):

or doesn't go that way ?

OpenStudy (amistre64):

row vectors are something different ... but yes, you could look at it like that

OpenStudy (amistre64):

spose we have a 1xn matrix; each column vector has 1 part; and there are n column vectors that span the space

OpenStudy (amistre64):

\[A = \{a_1,a_2,a_3,a_4,...,a_n\}\] each column vector in B has n parts \[B = \{b_1,b_2,b_3,b_4,...,b_p\}\] \[b_1=\begin{pmatrix}x_1\\x_2\\x_3\\x_4\\x_5\\.\\.\\.\\x_n\end{pmatrix}\]

OpenStudy (amistre64):

ive got no good way to try to define the AB operation off the fly

OpenStudy (christos):

I have already a definition of this thing here

OpenStudy (christos):

\[\sum\limits_{k = 1}^n {aikbkj} \]

OpenStudy (christos):

i and k are "smaller" letters down a and b

OpenStudy (christos):

but yet that definition I know it

OpenStudy (amistre64):

i = 1 by default since A only has 1 row

OpenStudy (christos):

ok

OpenStudy (amistre64):

\[\sum\limits_{k = 1}^n {a_{1k}~b_{kj}}~:~j=\{1,2,3,4,...,p\}\]

OpenStudy (amistre64):

im not sure about how i stated j tho

OpenStudy (christos):

hmm what do you mean by this part ? : j={1,2,3,4,...,p}

OpenStudy (christos):

ok ok I got what you mean

OpenStudy (christos):

thats the j of B

OpenStudy (amistre64):

yeah, since there are p column vectors in total, the inner product will contain 1 column with n parts; but each jth part is a calcuation of Arow1, and B column j

OpenStudy (christos):

do you base this on a being 1xp and also being the "first" in the inner product operation ? aB

OpenStudy (amistre64):

\[\begin{pmatrix}a_1~a_2~a_3\end{pmatrix}\begin{pmatrix}b_{11}~b_{12}\\b_{21}~b_{22}\\b_{31}~b_{32}\\\end{pmatrix}=\begin{pmatrix}a_1b_{11}+a_2b_{12}+a_3b_{13}\\a_1b_{21}+a_2b_{22}+a_3b_{23}\\a_1b_{31}+a_2b_{32}+a_3b_{33}\\\end{pmatrix}\]

OpenStudy (christos):

ok

OpenStudy (christos):

I see

OpenStudy (amistre64):

in trying to code that out, i did not produce the "second" column vector ....

OpenStudy (christos):

on which of the 3 matrices ?

OpenStudy (christos):

ah the third, its np

OpenStudy (amistre64):

\[\begin{pmatrix}a_1~a_2~a_3\end{pmatrix}\begin{pmatrix}b_{11}~b_{12}\\b_{21}~b_{22}\\b_{31}~b_{32}\\\end{pmatrix}=\begin{pmatrix}(a_1b_{11}+a_2b_{21}+a_3b_{31})~~(a_1b_{12}+a_2b_{22}+a_3b_{32}) \end{pmatrix}\]

OpenStudy (christos):

ok

OpenStudy (christos):

so apparently thats aB

OpenStudy (christos):

inner product

OpenStudy (amistre64):

AB should have 1 row, with p columns each column is a linear combination of the proper parts :)

OpenStudy (christos):

"proper parts" ?

OpenStudy (amistre64):

yeah, how to prove it tho ... im so far removed from any thrms and such to know what would make that more useful of simpler

OpenStudy (amistre64):

yes, an inner product is specifically defined; and the proper parts need to be used appropriately to fit the definition

OpenStudy (christos):

I can prove it just fine, the problem is

OpenStudy (christos):

that I don't quite understand what I have to prove :D

OpenStudy (christos):

I mean I understand kind of , but not entirely

OpenStudy (amistre64):

prove that AB is properly defined: A has n columns, B has n rows ..... they match and can be worked is the result a linear combination? that will have to be textbook defined to me

OpenStudy (christos):

so the aB has is a linear combination of the rows of B

OpenStudy (christos):

but how can it even be a linear combination of any sort since its a matrix * matrix = matrix (not linear combination)

OpenStudy (christos):

:D

OpenStudy (christos):

so basically I will find a 1xp AB matrix

OpenStudy (christos):

that's x + 2y + 3z ..... +tp (thats the linear combination ....? )

OpenStudy (amistre64):

if memory serves; a linear combination is defines as: For a given span: v1, v2, v3, ..., vk a linear combination is: av1 + bv2 + cv3 + ... + n vk

OpenStudy (christos):

and THOSE coefficients has to be the elements of a = { blah , blah , blah }

OpenStudy (christos):

so aB = a

OpenStudy (amistre64):

let A = {a1, a2, a3, ... , an} let B = {b1, b2, b3, ... , bp} let C = {c1,c2,c3,...,cp} show: c1 = a1b11 +a2b12 +a3b13 + ... + anb1n c2 = a1b21 +a2b22 +a3b23 + ... + anb2n c3 = a1b31 +a2b32 +a3b33 + ... + anb3n .... cp = a1bp1 +a2bp2 +a3bp3 + ... + anbpn can be written as a linear combination of the rows of B, where the coefficients are the entries of a

OpenStudy (amistre64):

class is starting .... gotta get